OpenStudy (anonymous):

evaluate where C is the line integral where C is the given curve:

OpenStudy (anonymous):

\[\int\limits xy dS\] C: x=t^2 y = 2t 0<t<1

OpenStudy (experimentx):

same procedure as before 1) change x and y to t's 2) \( ds = \sqrt{(\frac{dx}{dt})^2 + (\frac{dy}{dt})^2} dt\)